OpenStudy (otonogold):

Multiple Choice! Algebra 2! What is the number of real solutions? -11x^2=x+11 A) cannot be determined B) one solution C) two solutions D) no real solutions

OpenStudy (agent0smith):

First get it equal to zero (get everything on one side) Then find the discriminant.

OpenStudy (otonogold):

11x^2+x+11=0

OpenStudy (agent0smith):

For ax^2+bx+c = 0, the discriminant is b^2-4ac If it is positive, there are 2 real solutions If it's zero, there is 1 real solution If it's negative, there's 2 non-real complex solutions

OpenStudy (otonogold):

1 real solution?

OpenStudy (agent0smith):

\[\large 11x^2+x+11=0\] compare to\[\large ax^2 +bx+c=0\] find the value of \[\large b^2-4ac\]

OpenStudy (agent0smith):

What are a, b and c in your equation?

OpenStudy (otonogold):

-484

OpenStudy (agent0smith):

Did you forget the b^2? b is 1, a is 11, c is 11 so\[\large 1^2 - 4*11*11 = -483\]

OpenStudy (otonogold):

yes so no solution

OpenStudy (otonogold):

?

OpenStudy (agent0smith):

No real solutions, yes.

OpenStudy (otonogold):

Thank You

OpenStudy (agent0smith):

Welcome :)